Shift summary for day shift, 25/Nov/99

AG (
Thu, 25 Nov 1999 16:38:39 -0500 (EST)

Shift Summary for day shift, 25/Nov/99

On Shift: MA, AG

Hours of beam available at all intensities: 8.0
Hours of beam available at nominal intensity: 8.0
Hours of E799 data (special runs not included): 7.8
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

o stopping the run, changing tapes


Runs ended this shift:
Run 15053 - 13.0M - BEAM_799
(205k on-line 3pi0D)

Tapes brought to FCC this shift:
(including tapes for on-line split)

UP[A-C]18[1-3] and UPS061.


Accesses during shift:



Overview of shift:

Smooth E799 data-taking. Persisting SA4/SEM alarm forced us to change
limits from 1.1e-06 to 1.3e-6.


Description of Shift:

run spill time comment


15053_________0700_______Day shift begins. Run 15053 is in progress.
_________________________SA4/SEM ratio alarm goes on frequently, but
_________________________monitoring doesn't show any problems.

____160-170___0854_______Beam goes away and back on again.

______________1000_______Mounted UPE013.

_______321____1230_______Valerie changed RF by -1 ns. After a few krends
_________________________and pipe-line errors everything is back to normal

______________1330_______Tapes full - end run 15053 ...

______________1346_______Begin run 15054, but have donkeys all over the
_________________________place - do FIXDISP.

______________1330_______Changed limit SA4/SEM scaler alarm from 1.1_e6 to

______________1445_______trd_gas process in SDAQ was in error. Ashkan and
_________________________Theo (over the phone) had to reboot the serious
_________________________alarm terminal.


Follow up Notes on Previous Shift


Items to Follow up on Next Shift

Do drt_fresh_start during the next run start to restore monitoring